sue had 25 dollars.she spent y dollars on lunch.she now has $19.
Answer by zoomkaboom4(62)

25-y = 19
-25 -25
-1y = -6
/-1 /-1
y = $6
Any more questions? Message me!
-Megan (Zoomkaboom4)